# Muting Whisper: A Universal Acoustic Adversarial Attack on Speech Foundation Models

### v1

@@ Abstract
Recent developments in large speech foundation models like Whisper have led to their widespread use in many automatic speech recognition (ASR) applications. These systems incorporate ‘special tokens’ in their vocabulary, such as <endoftext>, to guide their language generation process. However, we demonstrate that these tokens can be exploited by adversarial attacks to manipulate the model’s behavior. We propose a simple yet effective method to learn a universal acoustic realization of Whisper’s <endoftext> token, which, when prepended to any speech signal, encourages the model to ignore the speech and only transcribe the special token, effectively ‘muting’ the model. Our experiments demonstrate that the same, universal 0.64-second adversarial audio segment can successfully mute a target Whisper ASR model for over 97% of speech samples. Moreover, we find that this universal adversarial audio segment often transfers to new datasets and tasks. Overall this work demonstrates the vulnerability of Whisper models to ‘muting’ adversarial attacks, where such attacks can pose both risks and potential benefits in real-world settings: for example the attack can be used to bypass speech moderation systems, or conversely the attack can also be used to protect private speech data.  [^1] The code is available at: https://github.com/rainavyas/prepend_acoustic_attack.
@@ 1 Introduction
Figure 1: Universal acoustic adversarial audio segment when prepended to any speech signal mutes the Automatic Speech Recognition (ASR) system, such that an empty transcription is generated. The <endoftext> token (EOT) is a special token in the Whisper vocabulary used to indicate the end of the generated transcription.
@@
Our experiments, conducted across eight different Whisper ASR models, demonstrate that the same universal 0.64-second adversarial audio segment can successfully ‘mute’ Whisper models for more than 97% of unseen speech samples. We further find that there is a surprising level of transferability of this universal adversarial audio segment to different speech domains (we consider four diverse datasets) and can even transfer to different tasks - the adversarial audio segment can ‘mute’ Whisper when used for speech translation as well as transcription.
Muting Whisper has significant implications in high stakes settings. Automatic speech recognition (ASR) systems play a crucial role in detecting and moderating harmful content such as hate speech (MacAvaney et al., 2019) in audio or video recordings (Wu and Bhandary, 2020). Muting Whisper poses a risk of circumventing this moderation process. Adversaries could exploit this vulnerability to release harmful content to the public audience without detection. Nevertheless, muting Whisper also has potential positive implications for speech privacy protection (Cheng et al., 2024). In contexts where speech recordings are transmitted over a network, malicious actors may attempt to extract private data through automated transcription. In such cases, our proposed method of muting Whisper could serve as a form of speech privacy protection, similar to a ‘jamming’ signal, where the speech remains unchanged for human listeners but ASR systems are unable to extract its content. On the whole, through this work we demonstrate both the risks and benefits of universal acoustic adversarial attacks aimed at muting Whisper.
@@ 2 Related Work
We extend the research on adversarial attacks for modern ASR systems such as Whisper, by outlining a method to develop a truly practical and effective adversarial attack with a real-world targeted objective. Specifically, this work makes the following contributions to adversarial attack literature on ASR systems:
  • We develop a short (0.64-second) adversarial audio segment that can be prepended to a speech signal. Existing research only considers superimposing the adversarial audio signal, which is not a practical setting for real-life attacks.
@@
3 Automatic Speech Recognition: Whisper
An Automatic Speech Recognition (ASR) system maps an speech/audio signal of $N$ frames , $\mathbf{x}=x_{1:N}$, to the text, $\mathbf{y}=y_{1:M}$ uttered in the speech signal - this is the transcription of the audio with $M$ words/tokens. Whisper’s encoder-decoder architecture, $\mathcal{F}(\cdot)$ with parameters $\theta$ auto-regressively predicts a vector representing the probability distribution over the vocabulary of tokens, $\mathcal{V}$, for the next token $y_{m}$, with the speech, $\mathbf{x}=x_{1:N}$ at the encoder input and the previously decoded tokens, $\mathbf{y}^{*}_{<m}$ at the decoder input,
@@ 3 Automatic Speech Recognition: Whisper
During the decoding process various special tokens are used by the Whisper model to guide the token generation. The first token (input to the decoder) is set as <|startoftranscript|>, followed by a token to indicate the language, for example <en> for English. As the Whisper model is trained to perform two different speech processing tasks (transcription and speech translation), the next token is used to indicate the task, e.g., <|transcribe|> or <|translate|>. Hence we define $\mathbf{y}^{*}_{0}=$ <|startoftranscript|> <lang tag> <|task tag|>. With this initialization, further tokens are generated auto-regressively from the vocabulary, $\mathcal{V}$ following Equation 1 and Equation 2. The auto-regressive decoding ends when the <|endoftext|> special token is predicted by the decoder.
4 Universal Acoustic Prepend Attack
@@ 4.2 Prepend Attack
To perturb a speech signal, $\mathbf{x}=x_{1:N}$, it is simplest to prepend a short, adversarial audio segment of $T$ frames, $\tilde{\mathbf{x}}=\tilde{x}_{1:T}$, such that the perturbed speech signal is $\tilde{\mathbf{x}}\oplus\mathbf{x}$, where $\oplus$ represents concatenation in the raw audio space. Then, given Whisper’s encoder-decoder model in Equation 1, the optimal adversarial audio segment, $\hat{\tilde{\mathbf{a}}}$, to ‘mute’ Whisper as per the adversarial objective, can be given as finding the adversarial audio segment that maximizes the probability of generating the <endoftext> special token (abbreviated to eot) as the first transcribed token,
@@ 4.4 Imperceptibility
For a truly practical adversarial attack, it is important for the adversarial audio segment generated to be sufficiently imperceptible such that it is not flagged as suspicious when prepended to natural speech signals. We achieve this imperceptibility in two dimensions. First, we ensure that the adversarial audio segment is extremely short such that there is little time for a human listener to detect the abnormal speech. We specifically limit the number of frames in the adversarial audio segment to $T=10,120$, which corresponds to 0.64-seconds of audio for a 16kHz sampling frequency. Next, we limit the ‘power’ of the adversarial audio segment, to ensure the amplitude of the adversarial audio segment is not significant relative to natural speech. To limit the power, we introduce a constraint in the optimization objective of Equation 4 that limits the amplitude of the adversarial audio,
@@
where $||\cdot||_{\infty}$ represents the l-infinity norm. By default we set $\epsilon=0.02$, as on the log-mel scale this empirically represents audio signals with power lower than typical human speech signals (refer to Figure 2). The l-infinity norm constraint is incorporated during gradient-based learning of the adversarial audio segment $\hat{\tilde{\mathbf{x}}}$, by clamping the values at $\epsilon$ [^2] For gradient-based learning we use ADAM (Kingma and Ba, 2015) optimization with clamping after each gradient update as in typical Projected Gradient Descent algorithms (Madry et al., 2019). Note that in practical settings it may be undesirable to have extremely low values for $\epsilon$, as the adversarial audio segment may then be contaminated by low-amplitude background noise.
5 Prepend Adversarial Attack Evaluation
@@ 5.1 Attack Performance Evaluation
For a learnt universal acoustic adversarial segment trained to maximize the probability of the Whisper model generating the <endoftext> special token as its first token for any speech signal, as per Equation 4, we can evaluate the performance of the adversarial attack by computing the percentage of unseen test speech signals, $\varnothing$, for which the attack is able to successfully ‘mute’ the Whisper model,
@@ 5.2 Adversarial Sensitivity Analysis
Beyond simply measuring the success of the universal acoustic adversarial attack in ‘muting’ an ASR system, it is meaningful to have a tool to analyze the mechanism of the attack that explains its success and lack of success for specific speech signals. One such tool is in the form of saliency analysis, which can indicate the sensitivity of the ASR system’s predictions to different parts of the input audio. The frames in the input audio that the transcription is most sensitive to are the parts of the audio that dominate the ASR system’s decisions and this can explain which part of the input the ASR system is attending to the most.
For an auto-regressive ASR model, $\mathcal{F}(\cdot)$ defined in Equation 1, we can define the $m$-th saliency of the universal adversarial audio segment, $\hat{\tilde{\mathbf{x}}}$, as the gradient of the $m$-th transcribed token, $\tilde{y}^{*}_{m}$, as
@@
As we are interested primarily in the first generated token, we set $m=1$ in our saliency analysis.
@@ 6.1 Experimental Setup
Results are reported across five diverse and popular speech recognition datasets: LibriSpeech (LBS) (Panayotov et al., 2015), TED-LIUM3 (TED) (Hernandez et al., 2018), MGB (Bell et al., 2015); Artie Bias (Artie) (Meyer et al., 2020) and Fleurs (Conneau et al., 2022). Details for each dataset are provided in Section A.1. The universal acoustic attack segment is learnt using the development split of the LibriSpeech dataset (for the task of transcription). The attack is then evaluated on the LibriSpeech test split and to measure the transferability of the attack it is also evaluated on the other datasets (TED-LIUM3, MGB and Artie Bias). The universal acoustic segment is evaluated for task transferability by also evaluating on speech transcription and speech translation tasks using the Fleurs dataset test splits.
@@
Experimental results are given for the family of Whisper ASR models Radford et al. (2023). Model details and their performance (Word Error Rate) on the datasets in this work have been provided for reference in Section A.2.
@@
The universal acoustic prepend attack segment is trained on the LibriSpeech development split. The attack segment is trained as per Equation 4, where it is prepended to speech samples in the raw audio space. The attack segment length is set to be 0.64 seconds and its maximum amplitude to $\epsilon=0.02$, to satisfy the constraint of Equation 5. Hyperparameter settings for learning the universal acoustic attack segment are given in Section A.3.
@@ 6.2 Results
The universal prepend attack segment is trained (on the LBS development split) to make the ASR model generate only an <endoftext> token, i.e. transcribe nothing. Evaluating on the LBS test-split, Table 1 gives the percentage of successful attacks, $\varnothing$ and the average sequence length of predicted transcriptions (asl) for the different target speech recognition models with the same (per model) trained 0.64-second universal acoustic adversarial segment prepended to every speech sample. A comparison is made to the no attack setting, where the speech samples are not modified in any manner. For every target Whisper model, the universal acoustic prepend attack is extremely successful in ensuring the model does not transcribe the speech signals, with the percentage of successful attacks increasing from more than 97% for the medium models to 99.9% for the tiny models. Similarly, in all cases the asl is brought to less than 1.0, whereas for the unattacked speech the transcriptions have nearly 18 words on average. We also compare to a random audio segment prepended to the speech samples and we find that this behaves identically to the no attack setting, i.e. a random attack cannot ‘mute’ Whisper. Overall, Table 1 shows that regardless of the model size, a short 0.64-second universal acoustic adversarial audio segment can be prepended (imperceptibly) to almost all speech signals to conceal the contents from Whisper speech recognition models.
@@
medium $\varnothing$(%$\uparrow$ 0.0 97.8
@@
Table 1: The percentage of successfully ‘muted’ speech samples, $\varnothing$, where the first generated token is <endoftext>, and the Average Sequence Length (asl) of transcriptions, for the LBS dataset. Results are presented for no attack, and for a trained (per model) universal acoustic adversarial attack, where the same universal adversarial segment is prepended to each speech sample.
Figure 2 gives the Mel-spectrogram of a random speech sample from the LBS test set with a 0.64-second universal acoustic adversarial segment prepended to the speech signal (learnt for the medium.en model). This validates that $\epsilon=0.02$ is an appropriate imperceptibility setting as it ensures that the power of the adversarial segment is always less than $\sim 1.50$dB, which is significantly lower than a typical human speech signal in the LBS dataset that can range from 1dB to more than 3.5dB. It is interesting to note that the acoustic adversarial segment covers the full range of frequencies relatively uniformly, which means it is likely to sound like static noise to a human listener.  [^3] We provide the universal acoustic adversarial attack segment audio files for each Whisper model at https://github.com/rainavyas/prepend_acoustic_attack/tree/main/audio_attack_segments
Figure 2: Mel spectrogram of universal acoustic segment (0.64s) prepended to a random speech sample from LBS dataset (truncated to a total length of 3s).
@@
Section 5.2 describes saliency as a tool to measure the sensitivity of the ASR model to the adversarial and the natural speech segments of the input audio. The average saliencies for the LBS dataset are given in Table 4, with a comparison for the successful attack samples and the failed attack samples. It is clear that a successful attack results in the ASR model being significantly more sensitive to the adversarial segment, and conversely more sensitive to the speech signal when the attack fails. This demonstrates that the universal acoustic attack is operating as intended, as a successful attack encourages the model to attend more to the acoustic realization of the <endoftext> special token (the adversarial audio segment).  [^4] Appendix D illustrates the frame-level saliency for example input audio signals. It is also interesting to note that for successful attack samples the saliency is significantly higher, suggesting that success of the adversarial attack is very dependent on the exact learnt universal adversarial segment.
@@
The universal attack segment has been trained on a specific domain of data (LBS data) and there is a risk that the attack may not necessarily transfer to different, distributionally shifted speech domains. Therefore, in this section we investigate the impact of transferring the 0.64-second universal acoustic adversarial segment to different unseen (during training of the attack) datasets, representing a diverse range of domain distributional shifts. Table 5 presents the results. For all models and datasets, the universal acoustic attack is able to continue muting the Whisper models for more than 90% of samples. Although this is slightly lower than 97% success rate for the in-domain LBS dataset, 90% is still a significant success rate, suggesting that the adversarial segment truly represents an acoustic realization of the <endoftext> token, which universally prevents the transcription of different speech domains.
@@
Table 5: Attack transferability across datasets: the percentage of successfully ‘muted’ speech samples, $\varnothing$, and the Average Sequence Length (asl) of generated transcripts with the universal acoustic adversarial attack learnt on LBS and evaluated on other datasets. Ref is the average reference transcription length.
Beyond transferability across data distributions, we also investigate how well the universal acoustic adversarial attacks transfer across different speech processing tasks. As the multilingual Whisper models can be instructed to perform transcription or speech translation, we evaluate how effective the adversarial segment (trained on Whisper for transcription) is in muting Whisper when used for speech translation. Table 6 presents attack results for speech translation from French (fr), German (de), Russian (ru) and Korean (ko) to English, from the Fleurs dataset. Two main trends can be identified. First, the attack transfers extremely well for the smaller Whisper models, with attack success rate greater than 94%, but for the larger models the success rate can drop to less than even 20%. Second, it appears that the ‘further’ the source language from English, the lower the success rate, e.g., the attack transfers better for French than Korean in general. As would be expected, this demonstrates that although the universal attack segment (the acoustic realization of <endoftext>) is extremely effective for muting Whisper for English transcription, its potency is increasingly limited as we transfer to a different speech task (translation) for distant languages.
@@
To ensure the universal acoustic adversarial segment is not easy for a human listener to detect, we set the maximum amplitude to $\epsilon=0.02$ and the length of the segment to 0.64-seconds. In this section we explore how much stricter these imperceptibility constraints can be made during the training of the universal acoustic attack segments. Figure 3 shows how the attack success percentage, $\varnothing$ (successfully mute Whisper) changes as the audio segment length is decreased from 0.64-seconds. The larger a model, the greater the decay in attack success. Further, the multi-lingual models tend to have a much greater decay than their English-only counterparts, with the attack success rate reaching near 0% for every multi-lingual model for a segment of 16-seconds. Figure 4 equivalently presents the impact of reducing the maximum amplitude, $\epsilon$. A similar trend (although less clear) arises where the larger and the multi-lingual variants of the models have a greater drop in success rate with a smaller $\epsilon$. The relative robustness of the multi-lingual and larger models in extremely constrained attack settings can perhaps be explained simply by the fact these models have been trained on more data and thus it is more difficult to find an adversarial acoustic segment that is a genuine universal realization of the <endoftext> special token.

### v2

@@ Abstract
Recent developments in large speech foundation models like Whisper have led to their widespread use in many automatic speech recognition (ASR) applications. These systems incorporate ‘special tokens’ in their vocabulary, such as <|endoftext|>, to guide their language generation process. However, we demonstrate that these tokens can be exploited by adversarial attacks to manipulate the model’s behavior. We propose a simple yet effective method to learn a universal acoustic realization of Whisper’s <|endoftext|> token, which, when prepended to any speech signal, encourages the model to ignore the speech and only transcribe the special token, effectively ‘muting’ the model. Our experiments demonstrate that the same, universal 0.64-second adversarial audio segment can successfully mute a target Whisper ASR model for over 97% of speech samples. Moreover, we find that this universal adversarial audio segment often transfers to new datasets and tasks. Overall this work demonstrates the vulnerability of Whisper models to ‘muting’ adversarial attacks, where such attacks can pose both risks and potential benefits in real-world settings: for example the attack can be used to bypass speech moderation systems, or conversely the attack can also be used to protect private speech data.  [^1] The code is available at: https://github.com/rainavyas/prepend_acoustic_attack.
@@ 1 Introduction
Figure 1: Universal adversarial audio segment when prepended to any speech signal mutes Whisper, such that an empty transcription is generated. The <endoftext> token (EOT) is a special token in the Whisper vocabulary used to indicate the end of the generated transcription.
@@
Our experiments, conducted across eight different Whisper ASR models, demonstrate that the same universal 0.64-second adversarial audio segment can successfully ‘mute’ Whisper models for more than 97% of unseen speech samples. We further find that there is a surprising level of transferability of this universal adversarial audio segment to different speech domains (we consider four diverse datasets) and can even transfer to different tasks - the adversarial audio segment can ‘mute’ Whisper when used for speech translation as well as transcription. Muting Whisper has significant implications in high stakes settings. Automatic speech recognition (ASR) systems play a crucial role in detecting and moderating harmful content such as hate speech (MacAvaney et al., 2019) in audio or video recordings (Wu and Bhandary, 2020). Muting Whisper poses a risk of circumventing this moderation process. Adversaries could exploit this vulnerability to release harmful content to the public audience without detection. Nevertheless, muting Whisper also has potential positive implications for speech privacy protection (Cheng et al., 2024). In contexts where speech recordings are transmitted over a network, malicious actors may attempt to extract private data through automated transcription. In such cases, our proposed method of muting Whisper could serve as a form of speech privacy protection, similar to a ‘jamming’ signal. Overall, this work demonstrates the vulnerability of Whisper models to muting adversarial attacks, which can have negative or positive implications.
@@ 2 Related Work
We extend the research on adversarial attacks for modern ASR systems such as Whisper, by outlining a method to develop a truly practical and effective adversarial attack with a real-world targeted objective. Specifically, this work makes the following contributions:
  • We develop a short (0.64-second) adversarial audio segment that can be prepended to a speech signal. Existing research tends to consider superimposing the adversarial audio signal, which is not a practical setting for real-world attacks.
@@
3 Speech Processing: Whisper
Continuous-time speech is sampled such that the audio can be represented as a sequence of samples, $\mathbf{x}=x_{1:N}$. An Automatic Speech Recognition (ASR) system maps this sampled speech/audio signal, $\mathbf{x}$, to the text, $\mathbf{y}=y_{1:M}$ uttered in the speech signal - this is the transcription of the audio with $M$ words/tokens. Whisper’s encoder-decoder architecture, $\mathcal{F}(\cdot)$ with parameters $\theta$ auto-regressively predicts a vector representing the probability distribution over the vocabulary of tokens, $\mathcal{V}$, for the next token $y_{m}$, with the speech, $\mathbf{x}=x_{1:N}$ at the encoder input and the previously decoded tokens, $\mathbf{y}^{*}_{<m}$ at the decoder input,
@@ 3 Speech Processing: Whisper
During the decoding process various special tokens are used by the Whisper model to guide the token generation. The first token (input to the decoder) is set as <|startoftranscript|>, followed by a token to indicate the language, for example <en> for English. As the Whisper model is trained to perform two different speech processing tasks (transcription and speech translation), the next token is used to indicate the task, e.g., <|transcribe|> or <|translate|>. Hence we define $\mathbf{y}^{*}_{0}=$ <|startoftranscript|> <lang tag> <|task tag|>  [^2] Note that for the English-only variant of Whisper models, $y^{*}_{0}=$ <|startoftranscript|>. With this initialization, further tokens are generated auto-regressively from the vocabulary, $\mathcal{V}$ following Equation 1 and Equation 2. The auto-regressive decoding ends when the <|endoftext|> special token is predicted.
4 Universal Prepend Attack
@@ 4.2 Prepend Attack
To perturb a speech signal, $\mathbf{x}=x_{1:N}$, it is simplest to prepend a short, adversarial audio segment of $T$ frames, $\tilde{\mathbf{x}}=\tilde{x}_{1:T}$, such that the perturbed speech signal is $\tilde{\mathbf{x}}\oplus\mathbf{x}$, where $\oplus$ represents concatenation in the raw audio space. Then, given Whisper’s encoder-decoder model in Equation 1, the optimal adversarial audio segment, $\hat{\tilde{\mathbf{x}}}$, to ‘mute’ Whisper as per the adversarial objective, can be given as finding the adversarial audio segment that maximizes the probability of generating the <|endoftext|> special token (abbreviated to eot) as the first transcribed token,
@@ 4.4 Imperceptibility
For a truly practical adversarial attack, it is important for the adversarial audio segment generated to be sufficiently imperceptible such that it is not flagged as suspicious when prepended to natural speech signals. We achieve this imperceptibility in two dimensions. First, we ensure that the adversarial audio segment is extremely short such that there is little time for a human listener to detect the abnormal speech. We specifically limit the number of frames in the adversarial audio segment to $T=10240$, which corresponds to 0.64-seconds of audio for a 16kHz sampling frequency. Next, we limit the ‘power’ of the adversarial audio segment, to ensure the amplitude of the adversarial audio segment is not significant relative to natural speech. To limit the power, we introduce a constraint in the optimization objective of Equation 4 that limits the amplitude of the adversarial audio,
@@
where $||\cdot||_{\infty}$ represents the l-infinity norm. By default we set $\epsilon=0.02$, as on the log-mel scale this empirically represents audio signals with power lower than typical human speech signals (refer to Figure 2). The l-infinity norm constraint is incorporated during gradient-based learning of the adversarial audio segment $\hat{\tilde{\mathbf{x}}}$, by clamping the values at $\epsilon$ [^3] Clamping after each gradient update is typical in Projected Gradient Descent (Madry et al., 2019). Note that in practical settings it may be undesirable to have extremely low values for $\epsilon$, as the adversarial audio segment may then be contaminated by low-amplitude background noise.
5 Muting Attack Evaluation
@@ 5.1 Attack Performance Evaluation
For a learnt universal acoustic adversarial segment trained to maximize the probability of the Whisper model generating the <|endoftext|> special token as its first token for any speech signal, as per Equation 4, we can evaluate the performance of the adversarial attack by computing the percentage of unseen test speech signals, $\varnothing$, for which the attack is able to successfully ‘mute’ the Whisper model,
@@ 5.2 Adversarial Sensitivity Analysis
Beyond simply measuring the success of the acoustic adversarial attack in ‘muting’ an ASR system, it is meaningful to analyze the mechanism of the attack that explains its success and lack of success for specific speech signals. We can analyze the saliency of the input audio to determine the sensitivity of the Whisper’s predictions to different parts of the input audio. The frames in the input audio that the transcription is most sensitive to are the parts of the audio that dominate Whisper’s decisions. For a model, $\mathcal{F}(\cdot)$ defined in Equation 1, we can define the $m$-th saliency of the universal adversarial audio segment, $\hat{\tilde{\mathbf{x}}}$, as the gradient of the $m$-th transcribed token, $\tilde{y}^{*}_{m}$,
@@
As we are interested primarily in the first generated token, we set $m=1$ in our analysis.
@@ 6.1 Experimental Setup
Results are reported across five diverse and popular speech recognition datasets: LibriSpeech (LBS) (Panayotov et al., 2015), TED-LIUM3 (TED) (Hernandez et al., 2018), MGB (Bell et al., 2015); Artie Bias (Artie) (Meyer et al., 2020) and Fleurs (Conneau et al., 2022). Details for each dataset are provided in Section A.1. The universal acoustic attack segment is learnt using the development split of the LBS dataset. The attack is then evaluated on the LBS test split and to measure the transferability of the attack it is also evaluated on the other datasets (TED-LIUM3, MGB and Artie Bias). The attack is evaluated for task transferability by also evaluating on speech transcription and speech translation tasks using the Fleurs dataset test splits.
@@
Experimental results are given for the family of Whisper ASR models Radford et al. (2023). Model details and their performance (Word Error Rate) on the datasets have been provided for reference in Appendix A.2.
@@
The universal acoustic prepend attack segment is trained on the LibriSpeech development split. The attack segment is trained as per Equation 4, where it is prepended to speech samples in the raw audio space. The attack segment length is set to be 0.64 seconds and its maximum amplitude to $\epsilon=0.02$, to satisfy the constraint of Equation 5. Further Hyperparameter settings for training the universal acoustic attack segment are given in Appendix A.3.
@@ 6.2 Results
The universal prepend attack segment is trained (on the LBS development split) to make the ASR model generate only an <|endoftext|> token, i.e. transcribe nothing. Evaluating on the LBS test-split, Table 1 gives the percentage of successful attacks, $\varnothing$ and the average sequence length of predicted transcriptions (asl) for the different target speech recognition models with the same (per model) trained 0.64-second universal acoustic adversarial segment prepended to every speech sample. A comparison is made to the no attack setting, where the speech samples are not modified in any manner. For every target Whisper model, the universal acoustic prepend attack is extremely successful in ensuring the model does not transcribe the speech signals, with the percentage of successful attacks increasing from more than 97% for the medium models to 99.9% for the tiny models. Similarly, in all cases the asl is brought to less than 1.0, whereas for the unattacked speech the transcriptions have nearly 18 words on average. We also compare to a random audio segment prepended to the speech samples and we find that this behaves identically to the no attack setting, i.e. a random attack cannot ‘mute’ Whisper. Overall, Table 1 shows that regardless of the model size, a short 0.64-second universal acoustic adversarial audio segment can be prepended (imperceptibly) to almost all speech signals to conceal the contents from Whisper speech recognition models.
@@
medium $\varnothing$ (%) $\uparrow$ 0.0 97.8
@@
Table 1: The percentage of successfully ‘muted’ speech samples, $\varnothing$, where the first generated token is <|endoftext|>, and the Average Sequence Length (asl) of transcriptions, for the LBS dataset. Results are presented for no attack, and for a trained (per model) universal acoustic adversarial attack, where the same universal adversarial segment is prepended to each speech sample.
Figure 2 gives the Mel-spectrogram of a random speech sample from the LBS test set with a 0.64-second universal acoustic adversarial segment prepended to the speech signal (learnt for the medium.en model). This validates that $\epsilon=0.02$ is an appropriate imperceptibility setting as it ensures that the power of the adversarial segment is always less than $\sim 1.50$dB, which is significantly lower than a typical human speech signal in the LBS dataset that can range from 1dB to more than 3.5dB. It is interesting to note that the acoustic adversarial segment covers the full range of frequencies relatively uniformly, which means it is likely to sound similar to static noise to a human listener.
Figure 2: Mel spectrogram of universal acoustic segment (0.64s) prepended to a (truncated) random speech sample from LBS dataset.
@@
Table 2: Average Sequence Length (asl) of generated transcripts for successful attack samples and failed attack samples. A successful sample is where the universal acoustic attack causes the Whisper model to generate a zero-length transcription.
@@
Table 3: Word Error Rate (WER) and breakdown (insertions, deletions and substitutions) between the transcript generated with no attack and the transcript generated with the universal attack, for the failed attack samples only. A failed sample is where the universal attack is unable to make Whisper generate a zero-length transcription.
@@
Section 5.2 describes saliency as a tool to measure the sensitivity of the ASR model to the adversarial and the natural speech segments of the input audio. The average saliencies for the LBS dataset are given in Table 4, with a comparison for the successful attack samples and the failed attack samples. It is clear that a successful attack results in the ASR model being significantly more sensitive to the adversarial segment, and conversely more sensitive to the speech signal when the attack fails. This demonstrates that the universal acoustic attack is operating as intended, as a successful attack encourages the model to attend more to the acoustic realization of the <|endoftext|> special token (the adversarial audio segment).  [^4] Appendix D illustrates the frame-level saliency. It is also interesting to note that for successful attack samples the saliency is significantly higher, suggesting that success of the adversarial attack is very dependent on the exact learnt universal adversarial segment.
@@
Table 4: Average saliency for the adversarial segment and speech segment (across LBS dataset) for successful and failed samples. A successful sample is where the universal attack causes Whisper to generate a zero-length transcription.
@@
The universal attack segment has been trained on a specific domain of data (LBS data) and there is a risk that the attack may not necessarily transfer to different, distributionally shifted speech domains. Therefore, in this section we investigate the impact of transferring the 0.64-second universal acoustic adversarial segment to different unseen (during training of the attack) datasets, representing a diverse range of domain distributional shifts. Table 5 presents the results. For all models and datasets, the universal acoustic attack is able to continue muting the Whisper models for more than 90% of samples. Although this is slightly lower than 97% success rate for the in-domain LBS dataset, 90% is still a significant success rate, suggesting that the adversarial segment truly represents an acoustic realization of the <|endoftext|> token, which universally prevents the transcription of different speech domains.
@@
Table 5: Attack transferability across datasets: the percentage of successfully ‘muted’ speech samples, $\varnothing$, and the Average Sequence Length (asl) of generated transcripts with the universal acoustic attack learnt on LBS and evaluated on other datasets. Ref is the average reference transcription length.
Beyond transferability across data distributions, we also investigate how well the universal acoustic adversarial attacks transfer across different speech processing tasks. As the multilingual Whisper models can be instructed to perform transcription or speech translation, we evaluate how effective the adversarial segment (trained on Whisper for transcription) is in muting Whisper when used for speech translation. Table 6 presents attack results for speech translation from French (fr), German (de), Russian (ru) and Korean (ko) to English, from the Fleurs dataset. Two main trends can be identified. First, the attack transfers extremely well for the smaller Whisper models, with attack success rate greater than 94%, but for the larger models the success rate can drop to less than even 20%. Second, it appears that the ‘further’ the source language from English, the lower the success rate, e.g., the attack transfers better for French than Korean in general.
@@
Next, we explore transferability of the attack across different Whisper models: this is explored analytically and empirically in Appendix C. The key finding is that certain attacks can be trained to transfer across models, but due to fundamental differences in the acoustic representation of the <|endoftext|> token for different models, it is unlikely a muting attack will naively transfer to unseen models.
@@
In this section we explore how much stricter imperceptibility constraints can be made during the training of the universal acoustic attack segments. Figure 3 shows how the attack success percentage, $\varnothing$ (successfully mute Whisper) changes as the audio segment length is decreased from 0.64-seconds. The larger a model, the greater the decay in attack success. Further, the multi-lingual models tend to have a much greater decay than their English-only counterparts, with the attack success rate reaching near 0% for every multi-lingual model for a segment of 16-seconds. Figure 4 equivalently presents the impact of reducing the maximum amplitude, $\epsilon$. A similar trend (although less clear) arises where the larger and the multi-lingual variants of the models have a greater drop in success rate with a smaller $\epsilon$. The relative robustness of the multi-lingual and larger models in extremely constrained attack settings can perhaps be explained simply by the fact these models have been trained on more data and thus it is more difficult to find a universal realization of the <|endoftext|> token.